OpenStudy (anonymous):

A right cylinder has a diameter of 30 mm and a height of 40 mm? What is the volume?

OpenStudy (anonymous):

28260 cubic mm

Parth (parthkohli):

The volume of a cylinder is \(\pi r^2h\). Here the radius is 15mm. \(\Large \color{Black}{\Rightarrow 3.14 \times 15 \times 15 \times40 }\) \(\Large \color{Black}{\Rightarrow 3.14 \times 225 \times 40 }\) \(\Large \color{Black}{\Rightarrow 3.14 \times 9000 }\) \(\Large \color{Black}{\Rightarrow 28260mm^3 }\)